For the following reaction:

2 A(g) ⇌ B(g) + C(g)

Given the following information, what is the equilibrium pressure of B in bar?

Initial pressure of A = 2.632 bar

Initial pressure of B = 0 bar

Initial pressure of C = 0 bar

K = 9.7 x 10-4


1
Expert's answer
2021-02-28T06:19:55-0500

Solution:

2A(g) ⇌ B(g) + C(g)

ICE Table:



The Kp expression for the reaction:

Kp = PB × PC / PA2


Kp = 9.7×10-4

PB = PC = 0.5P

PA = 2.632 - P

Hence,

9.7×10-4 = 0.5P × 0.5P / (2.632 - P)2

(2.632 - P)2 = 257.732 × P2

2.632 - P = 16.054 × P

2.632 = 17.054 × P

P = 0.154 (bar)


PB = 0.5P = 0.5 × 0.154 bar = 0.077 bar

PB = 0.077 bar


Answer: The equilibrium pressure of B is 0.077 bar.